What is the answer to: x+2/6=x-1/12
Answer by richwmiller(17219)

multiply both sides by 12
2*(x+2)=x-1
2x+4=x-1
x=-5
check
(-5+2)/6=(-5-1)/12
-3/6=-6/12
-1/2=-1/2
ok